HETAV PATEL

Grade 11,

If g1 g2 g3 are the accleration due to gravity on the earth surface ,on the top of the mountain and inside a mine then decreasing order is

If g1 g2 g3 are the accleration due to gravity on the earth surface ,on the top of the mountain and inside a mine then decreasing order is
 

Grade:12th pass

2 Answers

Vikas TU
4 years ago
Dear student 
we know on the surface of earth, it is
g1 =  GM / R^2
Say at that height h, the gravitational acceleration is g2.
So we can write, mg2 = GMm / (R+h)^2
=> g2 = GM/(R+h)^2
as altitude h increases, the value of acceleration due to gravity falls.
as depth h increases, the value of acceleration due to gravity falls.
and the order will be 
g1 > g3 > g2 
Hope this helps 
Good Luck 
Yashraj
3 years ago
going down or going up value of g decreases 
g1 >g3 >g 2
i think it is just formula based 
good luck
......................................